Two charges are located on the positive x-axis of a coordinate system. Charge π‘ž1 = 2 Γ— 10βˆ’9𝐢 is 2cm from the origin, and charge π‘ž2 = βˆ’3 Γ— 10βˆ’9𝐢 is 4cm from the origin. What is the total force exerted by these two charges on a charge π‘ž3 = 5 Γ— 10βˆ’9𝐢 located at the origin?


1
Expert's answer
2022-03-18T09:16:20-0400

The net force

"F=-F_{23}+F_{13}=-k\frac{|q_2||q_3|}{r_{23}^2}+k\frac{|q_1||q_3|}{r_{13}^2}"

"F=-9\times 10^9\times \frac{3\times 10^{-9}\times 5\times 10^{-9}} {0.04^2}\\
+9\times 10^9\times \frac{2\times 10^{-9}\times 5\times 10^{-9}} {0.02^2}=1.4\times 10^{-4}\:\rm N"